# The number of cultural traits, evolving genealogies, and the descendant process

### v1

@@ C.1 First moment of $L_{n}$
which holds for $a,b,m\in\mathbb{N}$ with $a\leq b$ (see Chapter 2 of Graham et al. 2003). In our case, set $a=N-M$ with $M=\min(n,N-m)$ and $b=N-1$. If $n\leq N-m$, then $(N-M)^{\underline{m+1}}=(N-n)^{\underline{m+1}}$. If, on the other hand, $n>N-m$, we have $(N-M)^{\underline{m+1}}=m^{\underline{m+1}}=0$, but also $(N-n)^{\underline{m+1}}=0$ because $m>N-n$. So, in any case, $(N-M)^{\underline{m+1}}=(N-n)^{\underline{m+1}}$, and
@@ C.2 Second moment of $L_{n}$
$\begin{split}\mathbb{E}\left[L_{n}^{2}\right]&=2\sum_{m_{1}=0}^{N-1}\sum_{\ell_{1}=1}^{\min(n,N-m_{1})}\frac{1}{u}\left(\frac{s}{Nu}\right)^{m_{1}}(N-\ell_{1})^{\underline{m_{1}}}\\ &\qquad\times\sum_{m_{2}=0}^{N-1}\frac{1}{u}\left(\frac{s}{Nu}\right)^{m_{2}}\frac{N^{\underline{m_{2}+1}}-\{N-(m_{1}+\ell_{1})\}^{\underline{m_{2}+1}}}{m_{2}+1}\\ &=2\sum_{m_{1}=0}^{N-1}\sum_{m_{2}=0}^{N-1}\left(\frac{1}{u}\right)^{2}\left(\frac{s}{Nu}\right)^{m_{1}+m_{2}}\\ &\qquad\times\sum_{\ell_{1}=1}^{\min(n,N-m_{1})}\frac{(N-\ell_{1})^{\underline{m_{1}}}N^{\underline{m_{2}+1}}-(N-\ell_{1})^{\underline{m_{1}}}\{N-(m_{1}+\ell_{1})\}^{\underline{m_{2}+1}}}{m_{2}+1}\\ &=2\sum_{m_{1}=0}^{N-1}\sum_{m_{2}=0}^{N-1}\left(\frac{1}{u}\right)^{2}\left(\frac{s}{Nu}\right)^{m_{1}+m_{2}}\sum_{\ell_{1}=1}^{\min(n,N-m_{1})}\frac{(N-\ell_{1})^{\underline{m_{1}}}N^{\underline{m_{2}+1}}-(N-\ell_{1})^{\underline{m_{1}+m_{2}+1}}}{m_{2}+1}\\ &=2\sum_{m_{1}=0}^{N-1}\sum_{m_{2}=0}^{N-1}\left(\frac{1}{u}\right)^{2}\left(\frac{s}{Nu}\right)^{m_{1}+m_{2}}\\ &\qquad\times\left[\frac{\{N^{\underline{m_{1}+1}}-(N-n)^{\underline{m_{1}+1}}\}N^{\underline{m_{2}+1}}}{(m_{1}+1)(m_{2}+1)}-\frac{N^{\underline{m_{1}+m_{2}+2}}-(N-n)^{\underline{m_{1}+m_{2}+2}}}{(m_{1}+m_{2}+2)(m_{2}+1)}\right].\\ \end{split}$ (C7)
@@ D Derivation of the popularity spectrum
Recall that ${P_{l}}$ is the expected number of traits carried by exactly $l$ individuals in an equilibrium population of size $N$. We calculate $(P_{l})_{l\in[N]}$ by adapting the method of Aoki (2018) (see also Strimling et al. (2009); Fogarty et al. (2015, 2017)) to continuous time as follows. Death occurs at rate $u$ to every individual, and when there are $l$ individuals carrying the trait (that is, the trait has popularity $l$), this trait loses a carrier at death rate $d_{l}=ul$, upon which its popularity decreases by one. Every individual that carries a trait with popularity $l$ transmits this trait to some individual that lacks the trait at rate $s(N-l)/N$, so the trait gains a carrier and its popularity increases by one; altogether, therefore, we have birth events at rate $b_{l}=sl(N-l)/N$ when in state $l\in[N]$. Innovation events produce new traits of popularity 1 at rate $N\mu$ without affecting any other trait. Therefore, the $P_{l}$ obey the following system of differential equations:
@@
[…]
@@
(where the first step comes from the normalisation of the hypergeometric distribution and the last step is true because $(N-l)^{\underline{n}}N^{\underline{l}}=N^{\underline{n+l}}=(N-n)^{\underline{l}}N^{\underline{n}}$)), (D4) and (D5) together give

### v2

@@ C.1 First moment of $L_{n}$
which holds for $a,b,m\in\mathbb{N}$ with $a\leq b$ (see Chapter 2 of Graham et al. 2003). In our case, set a=NMa=N-M with $M=\min(n,N-m)$ and b=N1b=N-1. If $n\leq N-m$, then $(N-M)^{\underline{m+1}}=(N-n)^{\underline{m+1}}$. If, on the other hand, n>Nmn>N-m, we have $(N-M)^{\underline{m+1}}=m^{\underline{m+1}}=0$, but also $(N-n)^{\underline{m+1}}=0$ because m>Nnm>N-n. So, in any case, $(N-M)^{\underline{m+1}}=(N-n)^{\underline{m+1}}$, and
@@ C.2 Second moment of $L_{n}$
[…]
@@ D Derivation of the popularity spectrum
Recall that $P_{l}$} is the expected number of traits carried by exactly $l$ individuals in an equilibrium population of size $N$. We calculate $(P_{l})_{l\in[N]}$ by adapting the method of Aoki (2018) (see also Strimling et al. (2009); Fogarty et al. (2015, 2017)) to continuous time as follows. Death occurs at rate $u$ to every individual, and when there are $l$ individuals carrying the trait (that is, the trait has popularity $l$), this trait loses a carrier at death rate $d_{l}=ul$, upon which its popularity decreases by one. Every individual that carries a trait with popularity $l$ transmits this trait to some individual that lacks the trait at rate s(Nl)/Ns(N-l)/N, so the trait gains a carrier and its popularity increases by one; altogether, therefore, we have birth events at rate $b_{l}=sl(N-l)/N$ when in state $l\in[N]$. Innovation events produce new traits of popularity 1 at rate $N\mu$ without affecting any other trait. Therefore, the $P_{l}$ obey the system of differential equations
@@
${P_{l}}=\frac{b_{1}b_{2}\cdots b_{l-1}}{d_{2}d_{3}\cdots d_{l}}{P_{1}}\hskip 30.0pt\text{for }\;2\leq l\leq N,\quad\text{complemented by}\quad P_{1}=\frac{N\mu}{u},$ (D2)
@@
(where the first step comes from the normalisation of the hypergeometric distribution and the last step is true because $(N-l)^{\underline{n}}N^{\underline{l}}=N^{\underline{n+l}}=(N-n)^{\underline{l}}N^{\underline{n}}$), (D4) and (D5) together give